Finding [OH-] which is 100 times greater than [H+]

The problem is for part (c) below.Thank you.
Calculate [H+] for each of the following solutions,and indicate whether the solution is acidic, basic, orneutral.(a) [OH -] = 0.00041 M
WebAssign will check your answer for the correct number of significant figures.correct check mark M correct check mark

(b) [OH -] = 2.5 10-9 M
WebAssign will check your answer for the correct number of significant figures.correct check mark M correct check mark

(c) a solution in which [OH -] is 100 times greater than[H+]
WebAssign will check your answer for the correct number of significant figures.wrong check mark
Your answer differs from the correct answerby orders of magnitude. M
Answer #1
If pH is 1-6.9 then it is acidIf pH is 7.1 - 14 then it is Basea)pOH = -log[OH]=-log(0.00041)=3.38pH = 14-pOH= 14-3.38= 10.612So it is Baseb) pOH = -log(2.5*10-9)=8.602pH = 14-pOH= 14-8.602= 5.3Hence it is acidicc) we know that[H+][OH-] =10-14here [OH-] is 100 times getter then[H+]so 100[H+][H+] =10-14∴[H+] =10-8Total concentration of [H+] in solution[H+] =10-7 + 10-8= 10-7 (1+0.1)=10-7 (1.1)pH = -log[H+]=-log10-7 (1.1)= 7-log(1.1)= 7-0.0413= 6.95Hence it is acidic
answered by: namisha